OpenStudy (anonymous):

Please help. These two points determine a line: (0,3) and (2,5). Which of these other points , (4,7) , (4,8) , (4,10) Are also on the line ?

OpenStudy (anonymous):

do you know parametric equations ?

OpenStudy (anonymous):

(4,7) lie on the line. the line equation is x-y+3=0.

OpenStudy (anonymous):

Show work

OpenStudy (anonymous):

woah, parametrics? slope=1, y-3=1(x-0) --> y=x+3, 7=?4+3 yes.

OpenStudy (anonymous):

What ?

OpenStudy (anonymous):

just get the equation of the line from the two points using the formula (y-y1)=m(x-x10 tke any one of the point as (x1,y1) m=(y1-y2)/(x1-x2) find m

OpenStudy (anonymous):

not really, x = 2t y = 3 + 2t , t is on real

OpenStudy (anonymous):

rise = 5-3=2, run = 2-0=2, slope = 2/2=1, point slope form: point (0,3) y-3=1(x-0) can be rearranged to y=x+3 then enter the three points to see if any of them make sense, and (4,7) does.

OpenStudy (amistre64):

f(x) = x+3 (x,f(x)) (4,f(4)) (4,7)